A sample of 1300 computer chips revealed that 18% of the chips do not fail in the first 1000 hours of their use. The company's promotional literature states that 20% of the chips do not fail in the first 1000 hours of their use. The quality control manager wants to test the claim that the actual percentage that do not fail is different from the stated percentage. Is there enough evidence at the 0.02 level to support the manager's claim?

Step 4 of 7 : Determine the P-value of the test statistic. Round your answer to four decimal places.
Answer #1

Solution :

Given that,

n = 1300

\hat p = 0.18

Hypothesis:

H0 : P = 0.20   

Ha : P\neq 0.20

Test Statistic:

Z =  \hat p - \hat{} \hat{}P / ( \sqrt ((P * (1 - P)) / n)

  Z = 0.18- 0.20 /  ( \sqrt (( 0.20 * 0.80) / 1300)

Z = -1.802

Z = -1.802

P-value = 0.0715

Decision Region : If P-value \leq   \alpha = 0.02 then reject Null hypothesis(H0)

P-value = 0.0715 >  \alpha = 0.02 , Fail to reject H0

Fail to Reject the null hypothesis. There is not enough evidence at the 0.02 level to support the manager's claim.